OpenStudy (anonymous):

Rationalise the denominator of 3/root 7

OpenStudy (***[isuru]***):

hi, \[\frac{ 3 }{ \sqrt{7} } = \frac{ 3 }{ \sqrt{7} } \times \frac{\sqrt{7} }{ \sqrt{7} }= \frac{ 3\sqrt{7} }{7 }\] hope this will help ya !!!

OpenStudy (anonymous):

Thx)

OpenStudy (***[isuru]***):

u r welcome!!